Given the equation y=-2x + 3, and the point (-5, 8), write an equation for a parallel line.

Answer by nerdybill(7384)

Given the equation y=-2x + 3, and the point (-5, 8), write an equation for a parallel line.
.
Slope of:
y=-2x + 3
is -2 (from "slope-intercept" form above)
.
Now, we have:
m = -2
(x1,y1) = (-5,8)
.
Plug above into "point-slope" form:
y - y1 = m(x - x1)
y - 8 = -2(x - (-5))
y - 8 = -2(x+5)
y - 8 = -2x-10
y = -2x-10+8
y = -2x-2 (answer in "slope-intercept" form)